# Simple Sublinear Algorithms for $(\Delta+1)$ Vertex Coloring

### v3

@@ Abstract
While being a quite natural and simple-to-state theorem, PST suffers from two drawbacks. Firstly, all its known proofs require technical arguments that rely on sophisticated graph decompositions and probabilistic arguments. Secondly, finding the coloring of the graph from the sampled lists in an efficient manner requires a considerably complicated algorithm.
@@ 1 Introduction
Assadi, Chen, and Khanna [ACK19] addressed this question by designing various sublinear algorithms for the $(\Delta+1)$ vertex coloring problem. These are algorithms whose resource requirements are significantly smaller than their input size. Some canonical examples include $(a)$ graph streaming algorithms [FKM+04] that process graphs by making one pass over their edges and using a limited space; $(b)$ sublinear time algorithms [CRT01] that process graphs by querying its adjacency list/matrix and using limited time; or $(c)$ Massively Parallel Computation (MPC) algorithms [KSV10] that process graphs in synchronous rounds and a distributed manner over machines with limited communication (see Section 2.1 for definitions).
The key contribution of [ACK19] was proving the following (purely combinatorial) theorem about $(\Delta+1)$ coloring, termed the palette sparsification theorem (PST).
Theorem 1 (Palette Sparsification Theorem (PST) [ACK19]).
@@
The authors in [ACK19] used PST to design sublinear algorithms in all the aforementioned models using the following “sparsifying” nature of this result: Firstly, it is easy to see that to color $G$ from the sampled lists, we can ignore all edges $(u,v)\in E$ where $L(u)\cap L(v)=\emptyset$ since they do not create any conflict under any possible coloring. Secondly, for any edge $(u,v)\in E$, the probability that it still remains conflicting—the probability that two independently chosen $O(\log{n})$-lists from $(\Delta+1)$ colors intersect—is $O(\log^{2}\!{(n)}/\Delta)$. Combining these with the upper bound of $n\Delta/2$ on the number of edges in $G$ implies that in expectation (and even with high probability) there are only $O(n\log^{2}\!{(n)})$ “important” edges that the algorithm should consider. The sublinear algorithms can now be obtained from this in a simple way [^4] For instance, for a streaming algorithm, first sample all the $O(n\log{n})$ colors, and during the stream whenever an edge arrives, see if it is an important edge and if so store it. At the end, use Theorem 1 to color the graph using these stored edges. This leads to a single-pass streaming algorithm with $O(n\log^{3}\!{(n)})$ bits of space for $(\Delta+1)$ coloring..
Since its introduction in [ACK19], PST and its variants and generalizations have been studied in sublinear algorithms [CFG+19, BCG20, AKM22, CGS22, ACGS23], distributed computing [FHM23, FGH+23, HKNT22], and discrete mathematics [AA20, ABD22, KK23, KK24, HK24, AK24].
Despite its long list of applications, and perhaps even due to its generality, PST suffers from two drawbacks. Firstly, the proof of PST is considerably complicated and technical; it goes through the so-called sparse-dense decomposition of graphs due to [Ree98] (and the variant proved in [ACK19] itself, building on [HSS16]), and then a detailed three-phase approach for coloring each part in the decomposition differently using various probabilistic and random graph theory arguments [^5] All known proofs of PST for $(\Delta+1)$ coloring (or so-called $(\deg+1)$-(list) coloring) [ACK19, AA20, HKNT22, KK23, FGH+24, AK24] follow this decomposition plus three-phase approach and it was even pointed out in [KK23] that: “something of this type […] seems more or less unavoidable”.. Secondly, PST, as stated, is a purely combinatorial theorem and not an algorithmic one, and thus to obtain the coloring of $G$, one needs to run a rather complicated and non-greedy post-processing algorithm to find the coloring of $G$ from the sampled lists (given also the decomposition of the graph; see [ACK19] for more details).
@@
The benefit of our APST compared to the original PST of [ACK19] (and all its other alternative variants in [AA20, HKNT22, KK23, FGH+24, AK24]) is twofold: it admits a significantly simpler proof, while also allowing for coloring from the sampled list using the standard greedy algorithm itself. At the same time, it is also weaker than the original PST on two fronts: it requires $O(\log^{2}\!{(n)})$ list sizes per vertex as opposed to $O(\log{n})$ (which is similar to [HKNT22, FGH+24]) but much more importantly, it allows the vertices to have much larger list sizes in the worst-case and only bounds the average list sizes (this is different than all prior work on PST, and was inspired by the recent work of [FM24] on the communication complexity of $(\Delta+1)$ coloring). We note that asymmetric list sizes are necessary if we would like to color the graph greedily from the sampled colors [^7] Consider coloring a $(\Delta+1)$-clique. When coloring the last vertex $v$ of the clique in the greedy algorithm, there is only one color that can be assigned to $v$ but to ensure this color is sampled by $v$ even with a constant probability, we need $L(v)$ to have size $\Omega(\Delta)$. Our APST addresses this by allowing vertices that are colored later in the greedy algorithm to also sample more colors, while earlier vertices can stick with smaller list sizes..
Finally, APST allows for recovering the same type of sublinear algorithms as in [ACK19] with only an extra $\mbox{\rm poly}\!\log\!{(n)}$ overhead in their resources. In particular, we obtain the following sublinear algorithms for $(\Delta+1)$ vertex coloring:
@@
The number of passes in our streaming algorithm is clearly optimal. The number of rounds in the MPC algorithm is also asymptotically optimal (in fact, it is only two rounds, and even one round assuming access to public randomness). It was also proven in [ACK19] that the runtime of the sublinear time algorithm and the space of the streaming algorithm are nearly optimal up to $\mbox{\rm poly}\!\log{(n)}$ factors (for the streaming algorithm, storing the coloring itself requires this much space anyway). These constitute the simplest known sublinear algorithms for $(\Delta+1)$ vertex coloring. [^9] To our knowledge, no other streaming nor sublinear time algorithms beside [ACK19] have been developed for this problem but for MPC algorithms, [CFG+19, CDP21] have subsequently presented other algorithms for this problem.
@@
Finally, in Appendix A, we also present an even simpler sublinear time algorithm for $(\Delta+1)$ coloring whose proof is inspired by our APST but does not directly uses it. While qualitatively weaker, the benefit of this algorithm is that it is entirely self-contained and only requires elementary probabilistic arguments (not even concentration inequalities), and thus can be even more “classroom friendly” than the algorithms obtained from our APST.
@@ 2.1 Sublinear Models Considered in this Paper
In this model, the input graph $G=(V,E)$ is presented to the algorithm as a stream of its edges ordered in some arbitrary manner. The algorithm makes a single pass (or sometimes multiple ones) over this stream while using $\widetilde{O}(n)$ memory and at the end of the stream, outputs a solution to the problem on the input graph $G$. See [FKM+04] for more details.
@@
In this model, the input graph $G=(V,E)$ is presented to the algorithm via query access to its adjacency list and matrix or alternatively speaking, using degree-, neighbor-, and pair-queries. The algorithm can make its queries adaptively or non-adaptively and then at the end outputs a solution to the problem on the graph $G$. See [CRT01] for more details.
@@
In this model, the input graph $G=(V,E)$ is originally partitioned across multiple machines, each with $\widetilde{O}(n)$ memory. Computation happens in synchronous rounds wherein each machine can send and receive $\widetilde{O}(n)$-size messages. After the last round, one designated machine outputs a solution to the problem on the graph $G$. See [KSV10] for more details.
@@ 3 Asymmetric Palette Sparsification
We present our APST in a more general form that allows for coloring vertices from arbitrary palettes of size proportional to their individual degrees—instead of the same original palette of $\left\{1,2,\ldots,\Delta+1\right\}$ described earlier— namely, the $(\deg+1)$-list coloring problem. We then obtain the APST for $(\Delta+1)$ coloring as a simple corollary of this result. We remark that a PST version of $(\deg+1)$-list coloring (by sampling $\mbox{\rm polylog}{(n)}$ colors per vertex in the worst-case) has been previously established in [HKNT22]; see also [AA20].
Theorem 2 (Asymmetric Palette Sparsification Theorem (for ($\deg+1$)-list coloring)).
@@
Theorem 2 follows immediately from Lemma 3.1 and Lemma 3.3 proven below.
@@
$\operatorname*{{\mathbb{E}}}\left[\ell(u)\cdot\ell(v)\right]\leqslant\sum_{i\neq j}\operatorname*{\textnormal{Pr}}(\pi(v)=i\,\wedge\,\pi(u)=j)\cdot\frac{(40\,n\ln{(n)})^{2}}{i\cdot j}=\frac{1600\,n^{2}\ln^{2}{(n)}}{{n\cdot(n-1)}}\cdot\sum_{i=1}^{n}\sum_{j=1}^{i-1}\frac{1}{i\cdot j}=O(\log^{4}(n)),$
[…]
@@
[…]
@@
Proof.
We condition on the choice of $\pi$ and by union bound obtain that with high probability, the complement of the event in Claim 3.2 holds for all vertices. For any vertex $v\in V$, we say a color in $S(v)$ is available to $v$ iff it is not assigned to any neighbor of $v$ by the time we process $v$ in the greedy algorithm. Let $a(v)$ denote the number of available colors and note that $a(v)\geqslant\deg^{<}_{\pi}(v)+1$.
In the following, we fix a vertex $v\in V$ and further condition on the randomness of $L(u)$ for all vertices $u$ with $\pi(u)>\pi(v)$. As we are using a fully deterministic greedy algorithm, the conditionings so far fix the set of available colors and the values of $a(v)$ and $\ell(v)$, but $L(v)$ is still a random $\ell(v)$-subset of the $\deg(v)+1$ colors in $S(v)$. If $\pi(v)\leqslant 40n\ln(n)/(\deg(v)+1)$, we have $\ell(v)=\deg(v)+1$ which means $L(v)=S(v)$ and thus there exists an available color in $L(v)$, proving the claim for this vertex.
For the remaining vertices with $\pi(v)>40n\ln(n)/(\deg(v)+1)$, we can apply Claim 3.2 to have,
@@
$\leqslant\exp\left(-\frac{\deg(v)\cdot\pi(v)}{2n}\cdot\frac{40n\ln{n}}{\pi(v)}\cdot\frac{1}{\deg(v)+1}\right)\leqslant n^{-5}.$
[…]
We obtain our APST for $(\Delta+1)$ coloring described earlier as a direct corollary of Theorem 2.
@@
[…]
@@ 4 Sublinear Algorithms from Asymmetric Palette Sparsification
We now show how our asymmetric palette sparsification theorem in Corollary 3.4 can be used to obtain sublinear algorithms for $(\Delta+1)$ vertex coloring. These algorithms are more or less identical to the (exponential-time) sublinear algorithms of [ACK19] from their original palette sparsification theorem and we claim no novelty in this part [^10] We do note that however, the time-efficient algorithms of [ACK19] are considerably more complex. They first need to find a so-called sparse-dense decomposition of the input graph via sublinear algorithms. Then, this decomposition is used to color the final graph from the sampled colors using an algorithmic version of the proof of their palette sparsification theorem which in particular requires a non-greedy and not-so-simple approach.. Instead, we merely point out how the “asymmetry” in list-sizes in Corollary 3.4 does not weaken the performance of the resulting sublinear algorithms beyond some $\mbox{\rm polylog}{(n)}$-factors, but instead leads to much simpler post-processing algorithms for finding the coloring of the graph from the sampled lists.
Theorem 3.
@@
As stated earlier, the proof of Theorem 3 follows the same exact strategy as the (exponential-time) sublinear algorithms of [ACK19]. To do so, we need the following definition.
@@
Claim 4.2.
The list of colors $\mathcal{L}$ consists of $O(n\log^{2}{n})$ colors with certainity and the expected number of edges in $G_{\mathcal{L}}=(V,E_{\mathcal{L}})$ is $\operatorname*{{\mathbb{E}}}\left|{E_{\mathcal{L}}}\right|=O(n\log^{4}{n})$.
Proof.
By the list sizes property of Corollary 3.4, we already know that $\mathcal{L}$ consists of $O(n\log^{2}{n})$ colors with certainity. For the second part, for any edge $(u,v)$, we have,
@@
We now prove Theorem 3 for each family of sublinear algorithms separately. In the following, we prove the resource guarantees of the algorithms only in expectation instead of a worst-case bound that happens with certainity. However, using the standard trick of running $O(\log{n})$ copies of the algorithm in parallel, terminating any copy that uses more than twice the expected resources, and returning the answer of any of the remaining ones, we obtain the desired algorithms in Theorem 3 as well (this reduction only increases space/query/memory of algorithms with an $O(\log{n})$ multiplicative factor and the error probability with a $1/\mbox{\rm poly}(n)$ additive factor).
@@
At the beginning of the stream, sample the colors $\mathcal{L}$ using Corollary 3.4 and during the stream, only store the edges that belong to $G_{\mathcal{L}}$. This is possible as we only need a random permutation and maximum degree $\Delta$ in order to provide list sizes of Corollary 3.4 and sample the colors $\mathcal{L}$. In the end, run the greedy algorithm on $G_{\mathcal{L}}$ and return the coloring. Corollary 3.4 ensures that with high probability $G$ is (list-)colorable from the sampled lists which leads to a $(\Delta+1)$ coloring of the entire graph. 4.1 ensures that we only need to work with $G_{\mathcal{L}}$ at the end of the stream and not all of $G$, and Claim 4.2 bounds the space of the algorithm with $\widetilde{O}(n)$ space in expectation.
We can also implement this algorithm in dynamic streams by recovering the conflict graph using a sparse recovery sketch instead of explicitly storing each of its edges in the stream. See [AGM12] for more on dynamic streams.
Finally, the knowledge of $\Delta$ is not necessary for this algorithm and can be removed using the same argument as in [ACK19].
@@
By the same argument as in our graph streaming algorithm, we can sample the colors $\mathcal{L}$ using Corollary 3.4 and query all pairs of vertices $u\neq v\in V$ where $L(u)\cap L(v)$ is non-empty to find the edges of $G_{\mathcal{L}}$. The same analysis as in Claim 4.2 applied to the ${n}\choose{2}$} vertex-pairs (instead of $\leqslant n\Delta/2$ edges), ensures that the expected number of queries is $\widetilde{O}(n^{2}/\Delta)$. We can then color $G_{\mathcal{L}}$ using the greedy algorithm in $\widetilde{O}(n)$ time. The correctness follows from Corollary 3.4 and 4.1 as before. This algorithm only requires adjacency matrix access to $G$ and we run it when $\Delta\geqslant\sqrt{n}$ to obtain an $\widetilde{O}(n\sqrt{n})$ time/query algorithm. When $\Delta\leqslant\sqrt{n}$, we instead run the standard greedy algorithm using $O(n\Delta)=\widetilde{O}(n\sqrt{n})$ time using the adjacency list access to $G$.
@@
The algorithm is almost identical to the semi-streaming one. Suppose we have access to public randomness. Then, we can sample the lists $\mathcal{L}$ publicly, and each machine that has an edge in $G_{\mathcal{L}}$ can send it to a designated machine. This way, a single machine receives all of $G_{\mathcal{L}}$ and can color it greedily. The correctness follows from Corollary 3.4 and 4.1 and the memory needed for this designated machine will be $\widetilde{O}(n)$ in expectation by Claim 4.2. Finally, we can remove the public randomness by having one machine do the sampling first on its own and share it with all the remaining machines in $O(1)$ rounds using the standard MPC primitives of search and sort.

### v4

@@ Abstract
While being a quite natural and simple-to-state theorem, PST suffers from two drawbacks. Firstly, all its known proofs require technical arguments that rely on sophisticated graph decompositions and probabilistic arguments. Secondly, finding the coloring of the graph from the sampled lists efficiently requires a considerably complicated algorithm.
@@ 1 Introduction
Assadi, Chen, and Khanna [undeff] addressed this question by designing various sublinear algorithms for the $(\Delta+1)$ vertex coloring problem. These are algorithms whose resource requirements are significantly smaller than their input size. Some canonical examples include $(a)$ graph streaming algorithms [undefp] that process graphs by making one pass over their edges and using a limited space; $(b)$ sublinear time algorithms [undefm] that process graphs by querying its adjacency list/matrix and using limited time; or $(c)$ Massively Parallel Computation (MPC) algorithms [undefac] that process graphs in synchronous rounds and a distributed manner over machines with limited communication (see Section 2.1 for definitions).
The key contribution of [undeff] was proving the following (purely combinatorial) theorem about $(\Delta+1)$ coloring, termed the palette sparsification theorem (PST).
Theorem 1.1 (Palette Sparsification Theorem (PST) [undeff]).
@@
The authors in [undeff] used PST to design sublinear algorithms in all the aforementioned models using the following “sparsifying” nature of this result: Firstly, it is easy to see that to color $G$ from the sampled lists, we can ignore all edges $(u,v)\in E$ where $L(u)\cap L(v)=\emptyset$ since they do not create any conflict under any possible coloring. Secondly, for any edge $(u,v)\in E$, the probability that it still remains conflicting—the probability that two independently chosen $O(\log{n})$-lists from $(\Delta+1)$ colors intersect—is $O(\log^{2}\!{(n)}/\Delta)$. Combining these with the upper bound of $n\Delta/2$ on the number of edges in $G$ implies that in expectation (and even with high probability) there are only $O(n\log^{2}\!{(n)})$ “important” edges that the algorithm should consider. The sublinear algorithms can now be obtained from this in a simple way [^1] For instance, for a streaming algorithm, first sample all the $O(n\log{n})$ colors, and during the stream whenever an edge arrives, see if it is an important edge and if so store it. At the end, use Theorem 1.1 to color the graph using these stored edges. This leads to a single-pass streaming algorithm with $O(n\log^{3}\!{(n)})$ bits of space for $(\Delta+1)$ coloring..
Since its introduction in [undeff], PST and its variants and generalizations have been studied in sublinear algorithms [undefl, undefj, undefg, undefk, undefe], distributed computing [undefr, undeft, undefw], and discrete mathematics [undefa, undefb, undefaa, undefab, undefy, undefc].
Despite its long list of applications, and perhaps even due to its generality, PST suffers from two drawbacks. Firstly, the proof of PST is considerably complicated and technical; it goes through the so-called sparse-dense decomposition of graphs due to [undefaf] (and the variant proved in [undeff] itself, building on [undefx]), and then a detailed three-phase approach for coloring each part in the decomposition differently using various probabilistic and random graph theory arguments [^2] All known proofs of PST for $(\Delta+1)$ coloring (or so-called $(\deg+1)$-(list) coloring) [undeff, undefa, undefw, undefaa, undefs, undefc] follow this decomposition plus three-phase approach and it was even pointed out in [undefaa] that: “something of this type […] seems more or less unavoidable”.. Secondly, PST, as stated, is a purely combinatorial theorem and not an algorithmic one, and thus to obtain the coloring of $G$, one needs to run a rather complicated and non-greedy post-processing algorithm to find the coloring of $G$ from the sampled lists (given also the decomposition of the graph; see [undeff] for more details).
@@
The benefit of our APST compared to the original PST of [undeff] (and all its other alternative variants in [undefa, undefw, undefaa, undefs, undefc]) is twofold: it admits a significantly simpler proof, while also allowing for coloring from the sampled list using the standard greedy algorithm itself. At the same time, it is also weaker than the original PST on two fronts: it requires $O(\log^{2}\!{(n)})$ list sizes per vertex as opposed to $O(\log{n})$ (which is similar to [undefw, undefs]) but much more importantly, it allows the vertices to have much larger list sizes in the worst-case and only bounds the average list sizes (this is different from all prior work on PST, and was inspired by the recent work of [undefv] on the communication complexity of $(\Delta+1)$ coloring). We note that asymmetric list sizes are necessary if we would like to color the graph greedily from the sampled colors [^4] Consider coloring a $(\Delta+1)$-clique. When coloring the last vertex $v$ of the clique in the greedy algorithm, there is only one color that can be assigned to $v$ but to ensure this color is sampled by $v$ even with a constant probability, we need $L(v)$ to have size $\Omega(\Delta)$. Our APST addresses this by allowing vertices that are colored later in the greedy algorithm to also sample more colors, while earlier vertices can stick with smaller list sizes..
Finally, APST allows for recovering the same type of sublinear algorithms as in [undeff] with only an extra $\mbox{\rm poly}\!\log\!{(n)}$ overhead in their resources. In particular, we obtain the following sublinear algorithms for $(\Delta+1)$ vertex coloring:
@@
The number of passes in our streaming algorithm is clearly optimal. The number of rounds in the MPC algorithm is also asymptotically optimal (in fact, it is only two rounds, and even one round assuming access to public randomness). It was also proven in [undeff] that the runtime of the sublinear time algorithm and the space of the streaming algorithm are nearly optimal up to $\mbox{\rm poly}\!\log{(n)}$ factors (for the streaming algorithm, storing the coloring itself requires this much space anyway). These constitute the simplest known sublinear algorithms for $(\Delta+1)$ vertex coloring. [^6] To our knowledge, no other streaming nor sublinear time algorithms beside [undeff] have been developed for this problem but for MPC algorithms, [undefl, undefn] have subsequently presented other algorithms for this problem.
@@
Finally, in Appendix A, we also present an even simpler sublinear time algorithm for $(\Delta+1)$ coloring whose proof is inspired by our APST but does not directly use it. While qualitatively weaker, the benefit of this algorithm is that it is entirely self-contained and only requires elementary probabilistic arguments (not even concentration inequalities), and thus can be even more “classroom friendly” than the algorithms obtained from our APST.
@@ 2.1 Sublinear Models Considered in this Paper
In this model, the input graph $G=(V,E)$ is presented to the algorithm as a stream of its edges ordered in some arbitrary manner. The algorithm makes a single pass (or sometimes multiple ones) over this stream while using $\widetilde{O}(n)$ memory and at the end of the stream, outputs a solution to the problem on the input graph $G$. See [undefp] for more details.
@@
In this model, the input graph $G=(V,E)$ is presented to the algorithm via query access to its adjacency list and matrix or alternatively speaking, using degree-, neighbor-, and pair-queries. The algorithm can make its queries adaptively or non-adaptively and then at the end outputs a solution to the problem on the graph $G$. See [undefm] for more details.
@@
In this model, the input graph $G=(V,E)$ is originally partitioned across multiple machines, each with $\widetilde{O}(n)$ memory. Computation happens in synchronous rounds wherein each machine can send and receive $\widetilde{O}(n)$-size messages. After the last round, one designated machine outputs a solution to the problem on the graph $G$. See [undefac] for more details.
@@ 3 Asymmetric Palette Sparsification
We present our APST in a more general form that allows for coloring vertices from arbitrary palettes of size proportional to their individual degrees—instead of the same original palette of $\left\{1,2,\ldots,\Delta+1\right\}$ described earlier— namely, the $(\deg+1)$-list coloring problem. We then obtain the APST for $(\Delta+1)$ coloring as a simple corollary of this result. Furthermore, we remark that a PST version of $(\deg+1)$-list coloring (by sampling $\mbox{\rm polylog}{(n)}$ colors per vertex in the worst-case) has been previously established in [undefw]; see also [undefa].
Theorem 3.1 (Asymmetric Palette Sparsification Theorem (for ($\deg+1$)-list coloring)).
@@
Theorem 3.1 follows immediately from Lemma 3.2 and Lemma 3.5 proven below.
@@
$\mathbb{E}\left[\ell(u)\cdot\ell(v)\right]$ $\leqslant\sum_{i\neq j}\textnormal{Pr}(\pi(v)=i\,\wedge\,\pi(u)=j)\cdot\frac{(40\,n\ln{(n)})^{2}}{i\cdot j}$
$=\frac{1600\,n^{2}\ln^{2}{(n)}}{{n\cdot(n-1)}}\cdot\sum_{i=1}^{n}\sum_{j=1}^{i-1}\frac{1}{i\cdot j}$
$=O(\log^{4}(n)),$
again, by the sum of the harmonic series. This concludes the proof.
@@
where in the second line we used the lower bound on $i=\pi(v)-1$ in the condition. Since $\deg^{<}_{\pi}(v)=X$ and the bound holds for all choices of $\pi(v)$, we can conclude the proof.
@@
Proof 3.6.
We condition on the choice of $\pi$ and by union bound obtain that with high probability, the complement of the event in Claim 1 holds for all vertices. For any vertex $v\in V$, we say a color in $S(v)$ is available to $v$ iff it is not assigned to any neighbor of $v$ by the time we process $v$ in the greedy algorithm. Let $a(v)$ denote the number of available colors and note that $a(v)\geqslant\deg^{<}_{\pi}(v)+1$.
In the following, we fix a vertex $v\in V$ and further condition on the randomness of $L(u)$ for all vertices $u$ with $\pi(u)>\pi(v)$. As we are using a fully deterministic greedy algorithm, the conditioning so far fix the set of available colors and the values of $a(v)$ and $\ell(v)$, but $L(v)$ is still a random $\ell(v)$-subset of the $\deg(v)+1$ colors in $S(v)$. If $\pi(v)\leqslant 40n\ln(n)/(\deg(v)+1)$, we have $\ell(v)=\deg(v)+1$ which means $L(v)=S(v)$ and thus there exists an available color in $L(v)$, proving the claim for this vertex.
For the remaining vertices with $\pi(v)>40n\ln(n)/(\deg(v)+1)$, we can apply Claim 1 to have,
@@
$\leqslant\exp\left(-\frac{\deg(v)\cdot\pi(v)}{2n}\cdot\frac{40n\ln{n}}{\pi(v)}\cdot\frac{1}{\deg(v)+1}\right)$
$\leqslant n^{-5}.$
Thus, with high probability, we can color $v$ from $L(v)$ in the greedy algorithm. Taking a union bound over all vertices concludes the proof.
We obtain our APST for $(\Delta+1)$ coloring described earlier as a direct corollary of Theorem 3.1.
@@
This way, we have that $L(v)$ is still a random $\ell(v)$-subset of $\Delta+1]$ as desired by Corollary 3.7, but now, we also have that conditioned on the choice of $S(v)$, $L^{\prime}(v)$ is a random $\ell^{\prime}(v)$-subset of $S(v)$ as in Theorem 3.1. Thus, we can apply Theorem 3.1 to $G$ and $\left\{S(v)\mid v\in V\right\}$ and argue that if we go over vertices of $G$ in the decreasing order of $\pi$, the greedy algorithm with high probability can color the graph from $L^{\prime}(v)\subseteq L(v)$ also, concluding the proof.
@@ 4 Sublinear Algorithms from Asymmetric Palette Sparsification
We now show how our asymmetric palette sparsification theorem in Section 3 can be used to obtain sublinear algorithms for $(\Delta+1)$ vertex coloring. These algorithms are more or less identical to the (exponential-time) sublinear algorithms of [undeff] from their original palette sparsification theorem and we claim no novelty in this part [^7] We do note that however, the time-efficient algorithms of [undeff] are considerably more complex. They first need to find a so-called sparse-dense decomposition of the input graph via sublinear algorithms. Then, this decomposition is used to color the final graph from the sampled colors using an algorithmic version of the proof of their palette sparsification theorem which in particular requires a non-greedy and not-so-simple approach.. Instead, we merely point out how the “asymmetry” in list-sizes in Corollary 3.7 does not weaken the performance of the resulting sublinear algorithms beyond some $\mbox{\rm polylog}{(n)}$-factors, but instead leads to much simpler post-processing algorithms for finding the coloring of the graph from the sampled lists.
Theorem 4.1.
@@
As stated earlier, the proof of Theorem 4.1 follows the same exact strategy as the (exponential-time) sublinear algorithms of [undeff]. To do so, we need the following definition.
@@
Claim 2.
The list of colors $\mathcal{L}$ consists of $O(n\log^{2}{n})$ colors with certainty and the expected number of edges in $G_{\mathcal{L}}=(V,E_{\mathcal{L}})$ is $\operatorname*{{\mathbb{E}}}\left|{E_{\mathcal{L}}}\right|=O(n\log^{4}{n})$.
Proof 4.3.
By the list sizes property of Corollary 3.7, we already know that $\mathcal{L}$ consists of $O(n\log^{2}{n})$ colors with certainty. For the second part, for any edge $(u,v)$, we have,
@@
We now prove Theorem 4.1 for each family of sublinear algorithms separately. In the following, we prove the resource guarantees of the algorithms only in expectation instead of a worst-case bound that happens with certainty. However, using the standard trick of running $O(\log{n})$ copies of the algorithm in parallel, terminating any copy that uses more than twice the expected resources, and returning the answer of any of the remaining ones, we obtain the desired algorithms in Theorem 4.1 as well (this reduction only increases space/query/memory of algorithms with an $O(\log{n})$ multiplicative factor and the error probability with a $1/\mbox{\rm poly}(n)$ additive factor).
@@
At the beginning of the stream, sample the colors $\mathcal{L}$ using Corollary 3.7 and during the stream, only store the edges that belong to $G_{\mathcal{L}}$. This is possible as we only need a random permutation and maximum degree $\Delta$ in order to provide list sizes of Corollary 3.7 and sample the colors $\mathcal{L}$. In the end, run the greedy algorithm on $G_{\mathcal{L}}$ and return the coloring. Corollary 3.7 ensures that with high probability $G$ is (list-)colorable from the sampled lists which leads to a $(\Delta+1)$ coloring of the entire graph. Section 4 ensures that we only need to work with $G_{\mathcal{L}}$ at the end of the stream and not all of $G$, and Claim 2 bounds the space of the algorithm with $\widetilde{O}(n)$ space in expectation.
We can also implement this algorithm in dynamic streams by recovering the conflict graph using a sparse recovery sketch instead of explicitly storing each of its edges in the stream. See [undef] for more on dynamic streams.
Finally, the knowledge of $\Delta$ is not necessary for this algorithm and can be removed using the same argument as in [undeff].
@@
By the same argument as in our graph streaming algorithm, we can sample the colors $\mathcal{L}$ using Corollary 3.7 and query all pairs of vertices $u\neq v\in V$ where $L(u)\cap L(v)$ is non-empty to find the edges of $G_{\mathcal{L}}$. The same analysis as in Claim 2 applied to the $\binom{n}{2}$ vertex-pairs (instead of $\leqslant n\Delta/2$ edges), ensures that the expected number of queries is $\widetilde{O}(n^{2}/\Delta)$. We can then color $G_{\mathcal{L}}$ using the greedy algorithm in $\widetilde{O}(n)$ time. The correctness follows from Corollary 3.7 and Section 4 as before. This algorithm only requires adjacency matrix access to $G$ and we run it when $\Delta\geqslant\sqrt{n}$ to obtain an $\widetilde{O}(n\sqrt{n})$ time/query algorithm. When $\Delta\leqslant\sqrt{n}$, we instead run the standard greedy algorithm using $O(n\Delta)=\widetilde{O}(n\sqrt{n})$ time using the adjacency list access to $G$.
@@
The algorithm is almost identical to the semi-streaming one. Suppose we have access to public randomness. Then, we can sample the lists $\mathcal{L}$ publicly, and each machine that has an edge in $G_{\mathcal{L}}$ can send it to a designated machine. This way, a single machine receives all of $G_{\mathcal{L}}$ and can color it greedily. The correctness follows from Corollary 3.7 and Section 4 and the memory needed for this designated machine will be $\widetilde{O}(n)$ in expectation by Claim 2. Finally, we can remove the public randomness by having one machine do the sampling first on its own and share it with all the remaining machines in $O(1)$ rounds using the standard MPC primitives of search and sort.